OpenStudy (anonymous):

Solve the quadratic equation y2 - 5y = -4 by factoring.

OpenStudy (anonymous):

y^2 - 5y + 4 = 0 when factoring in the form (x + a)(x + b) you know it'll be x^2 + ax + bx + ab so you want: ab = 4 and ax + bx = -5x -- which is (a + b) = -5 [CAN BE SOLVED BY SUBSTITUTION (lmk if you want details on that) OR CAN BE SOLVED BY TRIAL AND ERROR/LOGIC] I'll show you the logic since substitution sometimes takes longer and mental math is awesome :) : so to have ab = 4 your options are either a and b are both 2, or one of them is 4 and one of them is 1. (note since +4, either both positive or both negative) to have it correspond with a + b = - 5 . as you can see, a and b are -4 and -1 (in any order in this case; rarely the case) so factoring is (x - 1)(x - 4)